If it takes one person to mow a lawn in 6 hours, and another person 4 hours to mow the same yard how long will it take them to mow the lawn together?
A 15 inch tv has the width of 12 inches. What is the height?
A pro basketball player has a vertical leap of about 30 inches, what is his hang time? Use the hang-time function V=48T^2. His hang time is ? seconds?
: If it takes one person to mow a lawn in 6 hours, and another person 4 hours to mow the same yard how long will it take them to mow the lawn together?
A 15 inch tv has the width of 12 inches. What is the height?
A pro basketball player has a vertical leap of about 30 inches, what is his hang time? Use the hang-time function V=48T^2. His hang time is ? seconds?

Answer by solver91311(2197)

The first guy can mow a lawn in 6 hours, so he can mow 1/6 of a lawn in 1 hour, the second guy can mow 1/4 of a lawn in 1 hour. Together they can mow 1/6 + 1/4 of a lawn in 1 hour, or 2/12 + 3/12 = 5/12 of a lawn in 1 hour, so it should take 12/5 hours for both of them to mow the lawn. 12/5 = 2 and 2/5. 1/5 hour is 60/5 = 12 minutes, so 2/5 hour is 24 minutes. So: 2 hours, 24 minutes.

A 15in TV is measured on the diagonal. If the width is 12 inches, then use Pythagoras:

L = sqrt(15^2-12^2)

You do the arithmetic


If V = 48T^2 gives you vertical displacement in inches (and I think that is right, but you didn't say) simply substitute 30 for V in the formula and solve for T. Divide both sides by 48 and take the square root. Again, the arithmetic is up to you.